OpenStudy (anonymous):

Find the volume of the solid bounded below by the rectangle R : 1≤x≤e, 1≤y≤2 in the xy -plane and above by the graph of z=f(x,y)=ln(x)y .

OpenStudy (rational):

please show ur attempt

OpenStudy (anonymous):

\[\int\limits\limits_{1}^{2}\int\limits\limits_{1}^{e}\frac{ \ln(x) }{ y }dxdy=\int\limits\limits_{1}^{2}\frac{ e }{ y }\ln(e)-e-\frac{ 1 }{ y }\ln(y)+1dy\] \[\int\limits_{1}^{2}\frac{ e }{ y }-e+1=eln(2)-2e+2-eln(1)-e+1\] \[\approx0.165887556904675\]

OpenStudy (amistre64):

what is the integral of ln(x) ?

OpenStudy (anonymous):

xln(x)-x

OpenStudy (amistre64):

yeah, that is what i was trying to remember off the top of my head :)

OpenStudy (amistre64):

xlnx - x evaluated at 1 and e elne - e - (1ln1 - 1) e -e +1 = 1

OpenStudy (amistre64):

so determine the integral of 1/y from 1 to 2

OpenStudy (anonymous):

I got it right the answer is ln(2), thanks!

OpenStudy (amistre64):

\[\int\limits\limits_{1}^{2}\int\limits\limits_{1}^{e}\frac{ \ln(x) }{ y }dxdy=\int_{1}^{2}\frac 1y\left[\int\limits\limits_{1}^{e}ln(x) dx\right]dy\]

OpenStudy (amistre64):

ln2 is good for me as well :)